Consider the following. (Give your answers correct to two decimal places.)

(a) Calculate the value for the test statistic, χ2, for Ho: σ2 = 18.5, n = 19, s2 = 17.
χ2* =

(b) Calculate the value for the test statistic, χ2, for Ho: σ2 = 31.3, n = 14, s = 5.6.
χ2* =

(c) Calculate the value for the test statistic, χ2, for Ho: σ = 42.8, n = 14, s = 38.4.
χ2* =

(d) Calculate the value for the test statistic, χ2, for Ho: σ = 11.2, n = 23, s2 = 176.
χ2* =
Answer #1

Solution :

(a)

= (n - 1)s 2 /

= (19 - 1) * 17 / 18.5

= 16.54

(b)

= (14 - 1) 5.62 / 31.3 = 13.02

(c)

= (14 - 1) 38.42 / 42.82 = 10.46

(d)

= (23 - 1) * 176 / 11.22 = 30.87